You have a deck of cards. You pull one card. If it is a heart, you win $5. If it is a black card, you lose $15. If it is the ace of diamonds, you win $50. Otherwise, you win $1. What is the expected value for this game? Should you play this game?
Answer #1

P(Heart) = 13/52 = 1/4

P(Black card) = 26/52 = 1/2

P(Ace of diamonds) = 1/52

P(Anthing ese) = 1 - (1/4 + 1/2 + 1/52) = 3/13

Hence,

Expected value

= $5 * (1/4) - $15 * (1/2) + $50 * (1/52) + $1 * (3/13)

= $ -5.06

You should not play this game because expected value is negative.